OpenStudy (anonymous):

simplify sqrt 20x^4* sqrt 5x

OpenStudy (angela210793):

=2x^2sqr5*sqr5x=10x^2sqrx

OpenStudy (anonymous):

how did you do this

OpenStudy (angela210793):

sqr20=sqr(4*5)=2sqr5 sqrx^4=x^2 sqr5x=sqr5x multiply sqr5*sqr5x=sqr25x=5sqrx so in the ans u'll have : 10x^2sqrx